Please subscribe to the ACO-status mailing list for future ACO operational advisories and alerts.

The BSP2 dislodged from the bottom on 4/11/18, 00:53 UTC, moved to 25-30 mab and is floating with the current. CTD3, BPR pressure and fluorometer data may be affected.

New ACO instruments were installed during the ACO-3 cruise on 2014/11/01. Click here for ACO-3 videos.

ALOHA Cabled Observatory
Junction Box Engineering Data
JBox Engineering Data
Temperature Pressure 1 :
Temperature Pressure 2 :
Temperature :
Temperature Regenerator :
Pressure :
Voltage (12) :
Voltage (48) :
Voltage reference :
8.7
8
10.1
6.8
4.6
12
47.8
2.496
°C
°C
°C
°C
psi
V
V
V
Date: May 20, 2018    19:12:12 UTC